# with Controllable Camera and Object Motions

### v2

@@ Abstract
Controlling the movements of dynamic objects and the camera within generated videos is a meaningful yet challenging task. Due to the lack of datasets with comprehensive motion annotations, existing algorithms can not simultaneously control the motions of both camera and objects, resulting in limited controllability over generated contents. To address this issue and facilitate the research in this field, we introduce a Synthetic Dataset for Free-Form Motion Control (SynFMC). The proposed SynFMC dataset includes diverse objects and environments and covers various motion patterns according to specific rules, simulating common and complex real-world scenarios. The complete 6D pose information facilitates models learning to disentangle the motion effects from objects and the camera in a video. To validate the effectiveness and generalization of SynFMC, we further propose a method, Free-Form Motion Control (FMC). FMC enables independent or simultaneous control of object and camera movements, producing high-fidelity videos. Moreover, it is compatible with various personalized text-to-image (T2I) models for different content styles. Extensive experiments demonstrate that the proposed FMC outperforms previous methods across multiple scenarios.
Figure 1: Generation pipeline of videos in the proposed Synthetic Dataset for Free-Form Motion Control (SynFMC). The figure presents an example for generating synthetic video with three objects: (1) First, the HDRI environment map and objects matching the environment are selected as the assets. (2) Then, the motion types of objects and camera are selected for trajectory generation. (3) The center region shows the resulting 3D animation sequence used for rendering. The rendered video and annotations are demonstrated in the last row.
@@ 1 Introduction
Controlling motion dynamics in video generation has received increasing attention recently [10, 46, 45, 9, 61, 19, 52, 21, 18, 42, 16], as this capability enables better customization and plays a crucial role in many applications. For example, in filmmaking, directors meticulously choreograph the movements of both actors and the camera. Consequently, precise control over object and camera motions in video offers creative flexibility, enabling even non-professionals to create compelling visuals.
@@
Despite recent progress, challenges remain in this field. A key limitation is the lack of high-quality datasets with comprehensive motion annotations. For controlling object movement [46, 43, 54, 60, 23, 26, 4, 47, 20, 34], the motion is primarily annotated as the trajectory in image space [46, 54]. This annotation, however, intertwines the dynamics of both objects and the camera. For example, a rightward trajectory could represent either a stationary camera with a moving object or a static object with a left-moving camera. Besides, existing commonly used datasets [50, 63, 57] for learning camera motion mainly focus on scenes with minimal object dynamics. Some human-centric synthetic datasets [45, 3, 51, 55] provide ground truth for both human subjects and camera motions within a global coordinate system, yet exhibit limited motion diversity and category variety. Another significant challenge is the absence of methods capable of jointly controlling both object and camera movements. For example, approaches like Motion-Zero [4] only animate objects within a restricted range [17, 47, 23], while methods like CameraCtrl [10] exclusively target camera motion [14, 48, 1]. MotionCtrl [46] recently attempts to train separate modules for object and camera control in a two-stage process. However, without access to video data containing complete 6D pose annotations for both elements, it struggles to achieve reliable, synchronized control of objects and cameras within the same scene.
To enable free-form control of camera and object dynamics, a dataset with comprehensive pose annotations of both is essential. However, acquiring such data in real world is challenging and typically requires specialized equipment and expertise. To address this, we introduce a Synthetic dataset for Free-Form Motion Control (SynFMC). Designed with an emphasis on quality and diversity, the dataset includes a rich array of animated object assets across various categories, along with diverse HDRI (High Dynamic Range Imaging) environment maps. A rule-based generation strategy is implemented to create trajectories for both objects and the camera, as shown in Figure 1. This strategy encompasses basic patterns and simulates challenging cinematographic shots. To enhance realism, essential attributes for each object asset, such as living environment, types of speed and size, are annotated using a Multimodal Large Language Model (MLLM) [5] and manual labeling, facilitating the generation of plausible motion paths. The SynFMC dataset also provides detailed annotations, including 6D pose information of objects and camera, instance segmentation maps, depth maps, and comprehensive descriptions of content and motion, supporting a wide range of research fields.
To further validate the effectiveness of the proposed SynFMC dataset and address the limitations of previous methods, we propose a Free-Form Motion Control (FMC) method. FMC mainly includes two components: Camera Motion Controller (CMC) and Object Motion Controller (OMC). Unlike previous methods [54, 46], our approach leverages complete pose annotations of both camera and objects to disentangle global (camera) and local (object) dynamics. As shown in Figure 2, FMC is adaptable to various personalized Text-to-Image (T2I) models, producing high-fidelity results across a range of styles. In addition, FMC provides flexible user interfaces for motion control. Users can simultaneously input trajectories for objects and camera by simply drawing curves or by specifying motion types for each (as detailed in Sec. 3.3), which are used by a rule-based strategy to generate their trajectories accordingly.
@@
Table 1: Comparison of the proposed SynFMC with existing datasets. The object/camera motion pattern columns apply only to synthetic datasets. “Limited Movement” means that motion is restricted to a limited range. In addition to offering a rich variety of object categories, SynFMC outperforms in both motion pattern variety and and comprehensive pose annotations for camera and objects.
@@
BEDLAM [3]   10K Synthetic Human Limited Movement Static Ground Truth 3D Human Pose
SynBody [51]   27K Synthetic Human Limited Movement Static Ground Truth 3D Human Pose
HumanVid-Syn [45] 100K Synthetic Human Limited Movement Limited Patterns Ground Truth 3D Human Pose
SynFMC (ours)   26K Synthetic Common Diverse Patterns Diverse Patterns Ground Truth Object Pose
@@
  • The proposed FMC method enables joint or independent control of camera and object motions in video generation, achieving high-quality results across diverse scenes.
@@ 2 Related Work
Text-To-Video Generation. In text-to-video (T2V) models [9, 12, 36, 13, 22, 8, 35, 59], spatial modules determine the visual content while temporal modules control motion dynamics. However, lots of online videos are of low-quality characterized by motion blur and watermarks. Recently, Animatediff [9] incorporates the Domain Adapter into spatial modules to alleviate this issue. Moreover, it is highly compatible with other network modules [58, 53, 30], leading to the widespread usage within the community.
Dataset with Motion Annotations. For object control, most datasets [46, 54] focus on operation within the image space. However, the motions of camera and object are coupled in this space, while limiting the movement scope. On the other hand, only a few real datasets [50, 63, 57] contain camera pose information, and these primarily focus on static scenes without dynamic objects. Some synthetic datasets [51, 3, 55, 45] provide pose annotations for both objects and camera. Nevertheless, the 3D assets in these works are primarily focused on humans, resulting in limited category diversity. More discussions are in Sec 3.1.
Motion Control Methods. Most existing works [10, 17, 48, 60] can only control either object motion or viewpoint change. For methods that achieve manipulation over the motions of both the camera and objects, Direct-a-Video [50] can only simulate basic movements and simultaneous control in MotionCtrl [46] often results in suboptimal outcomes as noted in its study [46]. More discussions are in Sec 4.
@@ 3.1 Comparison with Existing Datasets
There is currently a lack of datasets that contain 6D poses for both objects and the camera [63, 57, 45]. As shown in Table 1, only a few real datasets [63, 57] provide estimated camera pose annotations, and they are primarily limited to scenes without dynamic objects due to the suboptimal performance of estimation methods [31, 37, 62]. To alleviate this, some methods [54, 46] use in-the-wild videos with image-space object trajectory inferred by optical flow models [62], but this way couples object and camera motions, resulting in ambiguities.
In contrast, synthetic datasets can conveniently obtain pose information [3, 55, 45, 51]. As shown in Table 1, most efforts focus on animating human. For example, HumanVid [45] constructs synthetic videos for controlling both human animation and camera motion. However, other common object categories remain underexplored. Furthermore, these datasets are limited to small human movements and simple camera motion patterns, restricting the model’s ability to learn complicated dynamic.
To effectively disentangle object and camera motion dynamics in video, it is essential to construct a new dataset with comprehensive object and camera poses. However, this is highly challenging in real world. First, capturing videos with complex, irregular object and camera motions is extremely difficult, typically requiring specialized equipment and expertise. Then, obtaining accurate pose estimation is difficult. Devices capable of capturing 6D poses for camera or objects are expensive and difficult to operate. Some studies [63, 46] attempt to obtain camera pose via estimation models. However, existing algorithms [62, 37] are time-intensive and often struggle with monocular videos containing dynamic objects. Besides, inferring 6D poses for general objects remains challenging. To address these limitations, we introduce SynFMC, a synthetic dataset generated using Unreal Engine [39], containing animations with diverse motion patterns and complete annotations.
@@ 3.2 Overview of SynFMC Dataset
SynFMC contains 26K videos divided into four groups: 6K static single-object, 6K static multi-object, 8K dynamic single-object, and 6K dynamic multi-object. Static means fixed object locations in world space while the camera remains movable. For diversity, the dataset includes common objects across various categories, such as humans, animals, plants and vehicles, as well as a wide range of environments like streets, grasslands, skies, oceans, etc. Additionally, SynFMC has diverse and complex multi-object and camera movements, covering not only basic motions but also shots that are challenging to achieve in real-world settings. Object assets are annotated with attributes such as speed and size by human annotators with the aid of MLLM [5] to ensure realistic motion simulation.
$\bullet$ Video Annotations. The proposed SynFMC dataset offers thorough annotations, including pose information for both objects and the camera, instance segmentation masks, depth maps, and detailed descriptions of content and motion, broadening its applicability across various research fields.
Figure 3: Object motion types. The line/curve in (c)/(d) are not limited in the horizontal plane. The trajectory of stationary point is not presented in the figure, which is a fixed point in the space.
@@ 3.3 Data Generation Pipeline
The key stages in the generation pipeline are outlined below.
$\bullet$ Asset Collection and Annotation. We collect a diverse set of 3D assets, including HDRI environment maps and animated objects, to enrich rendered content. The HDRI maps, used as backgrounds, span five types: groundnear groundskywater surface, and underwater, collected from PolyHaven [25]. For object assets, we select 6K resources with available animations from Objaverse-LVIS [7], Objaverse-XL [6], Mixamo [24], etc., covering a diverse range of categories. Low-quality assets are filtered out by human annotators. Then InternVL [5] is used to query basic properties of the object, such as class name, habitat, speed, and size category. This information is then verified or corrected by human annotators, who also label the motion type and provide a description for each animation.
$\bullet$ Object Motion. To create realistic motions, we design trajectories based on Bézier curves, where the rotations are derived from tangent and normal vectors along the curve. Examples of each motion type are shown in Figure 3. In implementation, the locations of control points are constrained within a reasonable range based on the object’s speed type.
$\bullet$ Camera Motion. For camera simulation, we decompose the relative movement of camera to object into 3 aspects: viewpoint type, distance type, and height type, see Figure 4. (a) Viewpoint type controls the orientation of the camera relative to the object: 1) front/back: camera captures objects from front/back view; 2) left/right side: camera captures objects from left/right side; 3) top: camera captures objects from top view. In implementation, viewpoint types are randomly assigned to selected key frames, with viewpoints for other frames determined through interpolation, enabling the camera’s range of motion to encompass all orientations around the object. (b) Distance type controls the horizontal distance between the camera and objects: 1) zoom in/out and 2) static. (c) Height type controls the vertical distance between the camera and objects: 1) up/down and 2) static. To keep the object within the camera’s view without anchoring it at the image center, the camera orientation is adjusted to focus on a random offset from the object’s centroid.
@@
Figure 4: Decomposition of camera motion. We decompose the camera motion into three aspects. (a) Viewpoint type indicates the orientation of the camera relative to the object. 1 - 5 present front/back, left/right side and top perspectives. (b) Distance type and (c) Height type determine horizontal and vertical distance between the camera and the object respectively. 6 - 9 are zoom in/out and up/down respectively. The “static” types are omitted in (b) and (c), which stand for fixed distances.
Table 2: Comparison of the proposed FMC with existing methods. FMC excels in more flexible inputs and a wide range of motion patterns.
@@
$\bullet$ Generation of Multi-Object Scenes. First, objects from the same environment with comparable sizes and speeds are selected. Next, we generate trajectories for each object. Specifically, the first object’s trajectory is created using the way described earlier, while the trajectories of others are generated based on the previous object’s path, following a similar way used in camera motion simulation. For camera motion, we randomly select an object as the tracking target in each segment, with the trajectory generated using the same method as previously outlined.
$\bullet$ Rendering. This stage creates synthetic videos based on selected assets and motion types of objects and the camera as shown in Figure 1. Specifically, for each segment in the video, we randomly select animations for the chosen objects and generate trajectories according to its annotated motion type. Similarly, random combinations of camera motion types are applied across different segments, allowing the generated videos to encompass a wide range of motion patterns as in real-world scenarios. Finally, the selected assets (background and objects) and poses of the camera and objects are imported into Unreal Engine [39] to get 3D animation sequence for rendering.

### v3

# and Objects in Video Generation

@@ Abstract
Controlling the movements of dynamic objects and the camera within generated videos is a meaningful yet challenging task. Due to the lack of datasets with comprehensive 6D pose annotations, existing text-to-video methods can not simultaneously control the motions of both camera and objects in 3D-aware manner, resulting in limited controllability over generated contents. To address this issue and facilitate the research in this field, we introduce a Synthetic Dataset for Free-Form Motion Control (SynFMC). The proposed SynFMC dataset includes diverse object and environment categories and covers various motion patterns according to specific rules, simulating common and complex real-world scenarios. The complete 6D pose information facilitates models learning to disentangle the motion effects from objects and the camera in a video. To provide precise 3D-aware motion control, we further propose a method trained on SynFMC, Free-Form Motion Control (FMC). FMC can control the 6D poses of objects and camera independently or simultaneously, producing high-fidelity videos. Moreover, it is compatible with various personalized text-to-image (T2I) models for different content styles. Extensive experiments demonstrate that the proposed FMC outperforms previous methods across multiple scenarios.
Figure 1: The rule-based generation pipeline of videos in the proposed Synthetic Dataset for Free-Form Motion Control (SynFMC). This example generates synthetic video with three objects: (1) The environment asset and it’s matching object assets are selected as the scene elements. (2) The motion types of objects and camera are randomly selected for trajectory generation. (3) The center region shows the resulting 3D animation sequence used for rendering. The rendered video and annotations are demonstrated in the last row.
@@ 1 Introduction
Controlling motion dynamics in video generation has received increasing attention [12, 47, 46, 11, 21, 53, 24, 20, 42, 18, 35], as it enables better customization and is crucial in many applications. For example, in filmmaking, directors meticulously choreograph the movements of both actors and the camera. Consequently, precise control over object and camera motions in video offers creative flexibility.
@@
Despite recent progress, challenges remain in motion control of text-to-video (T2V) generation. A key limitation is the lack of high-quality datasets with comprehensive 6D pose annotations. For controlling object movement [47, 23, 44, 54, 59, 25, 27, 4, 48, 22, 34], the motion is primarily annotated as the trajectory in image space [47, 54]. This annotation, however, lacks 3D nature and intertwines the dynamics of both objects and the camera. For example, a rightward trajectory could represent either a stationary camera with a moving object or a static object with a left-moving camera. Recently, 360°360°360  °-Motion synthetic dataset [10] provides 6D poses of objects, but limited with static camera setting and motion diversity. Besides, existing commonly used datasets [51, 61, 57] for learning camera motion mainly focus on scenes with minimal object dynamics. Some human-centric synthetic datasets [46, 3, 52, 55] provide ground truth for both human subjects and camera motions within a global coordinate system, yet exhibit limited motion diversity and category variety.
Another limitation is the absence of methods that can independently or jointly control the 6D poses of both object and camera. For example, methods like Motion-Zero [4] animate objects without 3D-aware control (e.g., orientation) [19, 48, 25], while methods like CameraCtrl [12] exclusively focus on camera motion [16, 49, 1]. MotionCtrl [47] trains separate modules for object and camera control in a two-stage process. However, without access to video data containing complete 6D pose annotations for both elements, it struggles to achieve realistic, synchronized control of objects and cameras within the same scene.
@@
To address these limitations, a dataset with comprehensive 6D pose annotations of objects and camera is desired. However, acquiring such data is challenging and typically requires specialized equipment and expertise. In this work, we introduce a Synthetic dataset for Free-Form Motion Control (SynFMC). Designed with an emphasis on quality and diversity, the dataset includes a rich array of animated object assets and environment assets across various categories. Furthermore, a rule-based generation algorithm is implemented to create trajectories for both objects and the camera, as shown in Fig. 1. This algorithm encompasses basic patterns and simulates challenging cinematographic shots as in Fig. 5. Compared to recent works [10, 46] that can only construct uncontrolled trajectories, our rule-based algorithm supports customized object & camera motions with diverse patterns. To enhance realism, essential attributes of objects, like living environment, types of speed and size, are annotated using a Multimodal Large Language Model (MLLM) [5] and manual labeling, facilitating the generation of plausible trajectories. The SynFMC dataset also provides detailed annotations, including 6D pose information of objects and camera, instance segmentation maps, depth maps, and comprehensive descriptions of content and motion, supporting a wide range of research fields.
To further validate the effectiveness of the proposed SynFMC dataset and support 3D-aware control in T2V generation, we propose a Free-Form Motion Control (FMC) method. FMC mainly includes two components: Camera Motion Controller (CMC) and Object Motion Controller (OMC). Unlike previous methods [54, 47], Our approach trained on SynFMC disentangles global (camera) and local (object) dynamics and manipulates the 6D poses of camera and objects. Furthermore, we adopt Domain LoRA [17] to prevent model from fitting to rendered style in synthetic data. As shown in Fig. 2 and Fig. 6, FMC effectively mitigates the domain gap and adapts to various personalized Text-to-Image (T2I) models, generating high-fidelity results across diverse styles. In addition, FMC provides flexible user interfaces for motion control. Users can input trajectories for objects and camera by simply drawing 3D curves or by specifying motion types for each (as detailed in Sec. 3.3), which are used by the rule-based algorithm to generate their trajectories accordingly.
Table 1: Comparison of the proposed SynFMC with existing datasets. The object/camera motion pattern columns apply only to synthetic datasets. In addition to offering a rich variety of object categories, SynFMC outperforms in motion pattern variety and controllability with comprehensive pose annotations of camera and objects. In our implementation, we only use 26K subset as training data.
@@
BEDLAM [3]   10K Synthetic Human Limited/Uncontrollable Static Ground Truth 3D Human Pose
SynBody [52]   27K Synthetic Human Limited/Uncontrollable Static Ground Truth 3D Human Pose
HumanVid-Syn [46] 100K Synthetic Human Limited/Uncontrollable Diverse/Uncontrollable Ground Truth 3D Human Pose
360°360°360  °-Motion [10]   54K Synthetic Animal & Human Limited/Uncontrollable Static Object Pose
SynFMC (ours)   62K Synthetic Common Diverse/Controllable Diverse/Controllable Ground Truth Object Pose
@@
  • The FMC method can manipulate the 6D poses of the camera and objects independently or simultaneously, achieving high-quality results across diverse scenes.
@@ 2 Related Work
Dataset with Motion Annotations. Most datasets [47, 54] focus on operation within the image space. However, camera and object motions are coupled in this space, while limiting the movement scope. On the other hand, only a few real datasets [51, 61, 57] provide camera pose, and these primarily focus on static scenes without dynamic objects. Some synthetic datasets [52, 3, 55, 46] provide pose annotations for both objects and camera, yet their 3D assets are predominantly human-centric, limiting category diversity. More discussions are in Sec 3.1.
Motion Control Methods. Most existing works [12, 19, 49, 59, 10] can only control either object motion or viewpoint change. For methods that support manipulation of both the camera and objects, Direct-a-Video [51] can only simulate basic movements and simultaneous control in MotionCtrl [47] often results in suboptimal outcomes as noted in its study [47]. More discussions are in Sec. 4.
@@ 3.1 Comparison with Existing Datasets
There is currently a lack of datasets that contain 6D poses of both objects and the camera [61, 57, 46]. As shown in Tab. 1, only a few real datasets [61, 57] provide estimated camera poses and are primarily limited to scenes without dynamic objects due to the suboptimal performance of estimation methods [31, 37, 60]. Some methods [54, 47] use in-the-wild videos with image-space object trajectory inferred by optical flow models [60], but this entangles object and camera motions while lacking 3D information like orientation. Synthetic datasets conveniently obtain pose information [3, 55, 46, 52], but most focus on human and are limited to small movements and simple camera motion patterns, limiting the ability to learn complex dynamic.
To facilitate model learning to control the motions in 3D-aware manner, it is essential to construct a new dataset with comprehensive object and camera poses. However, this is highly challenging in real world. First, capturing videos with complex, irregular object and camera motions is extremely difficult, typically requiring specialized equipment and expertise. Then, obtaining accurate pose estimation is difficult. Devices capable of capturing 6D poses for camera or objects are expensive and difficult to operate. Some studies [61, 47] attempt to obtain camera pose via estimation models. However, existing algorithms [60, 37] are time-intensive and often struggle with monocular videos containing dynamic objects. Besides, inferring 6D poses for general objects remains challenging. To address these limitations, we introduce SynFMC, a synthetic dataset generated using Unreal Engine [39], containing animations with diverse motion patterns and complete annotations.
$\bullet$ Difference from 360°360°360  °-Motion Dataset [10]. 1) We handle both static&dynamic cameras, enabling more complex shots than [10]’s static setup. 2) Our rule-based algorithm supports diverse (non-)horizontal object motions, unlike [10]’s GPT-derived horizontal-only ones. 3) Our environment & object assets extend beyond [10]’s terrestrial scenes.
$\bullet$ Difference from HumanVid-Syn Dataset [46]. 1) The object trajectories of [46] rely on predefined 3D motion assets (SMPL-X/skeleton), whereas our rule-based algorithm enables diverse patterns, e.g., in-place/(non-)horizontal motions in Fig. 32) [46] randomly samples camera positions within a semi-cylinder in front of human, while we achieve fine-grained control by decoupling camera motion (Fig. 4), supporting controllable and diverse movements. 3) [46] focuses on synthesizing single-human animation, whereas we support multi-object scenarios, generating richer dynamics.
@@ 3.2 Overview of SynFMC Dataset
SynFMC contains 62K videos divided into four groups: 15K static single-object, 15K static multi-object, 16K dynamic single-object, and 16K dynamic multi-object. Static means fixed object locations in world space while the camera remains movable. For diversity, the dataset includes common objects across various categories, such as humans, animals, plants and vehicles, as well as a wide range of environments like streets, grasslands, skies, oceans, etc. Additionally, SynFMC has diverse and complex multi-object and camera movements, covering not only basic motions but also shots that are challenging to achieve in real-world settings. Object assets are annotated with attributes such as speed and size by human annotators with the aid of MLLM [5] to ensure realistic motion simulation.
$\bullet$ Video Annotations. The proposed SynFMC dataset offers thorough annotations, including pose information for both objects and the camera, instance segmentation masks, depth maps, and detailed descriptions of content and motion, broadening its applicability across various fields [9, 8, 43].
Figure 3: Object motion types. The trajectory of stationary point is not presented in the figure, which is a fixed point in the space.
@@ 3.3 Data Generation Pipeline
$\bullet$ Asset Collection and Annotation. We collect diverse 3D assets, including environmens and objects. Environment assets span five types: groundnear groundskywater surface, and underwater. We collect high quality panoramic HDRI images from internet. For object assets, we select animated object assets from Objaverse-LVIS [7], Objaverse-XL [6], Mixamo [26], etc., covering diverse categories. Human annotators filter low-quality assets and verify/correct object properties (e.g., class, habitat, speed, size) queried by InternVL [5]. They also label motion type and provide description for each animation.
$\bullet$ Object Motion. To create realistic motions, we design trajectories based on Bézier curves in each motion segment, with rotations derived from tangent and normal vectors along the curve. Fig. 3 exemplifies motion types. Control points are constrained based on the object’s speed.
$\bullet$ Camera Motion. We decompose camera motion into 3 types: viewpointdistance, and height, see Fig. 4. (a) View-point controls camera’s orientation when capturing objects: front/back, left/right, and top. Start and end viewpoints of a motion segment are randomly assigned, with intermediate frames interpolated for smooth transitions, enabling camera’s motion range to encompass all orientations around the object. (b) Distance controls the horizontal distance between the camera and objects: zoom in/out and static. (c) Height controls the vertical distance: up/down and static. To maintain object visibility without centering it, the camera targets a randomly offset point near the object’s centroid.
@@
Figure 4: Camera motion types. We decompose camera motion into 3 aspects. (a) Viewpoint controls camera orientation when capturing object. 1 - 5 present front/back, left/right and top perspectives. (b) Distance and (c) Height determine horizontal and vertical distance between camera and object, respectively. 6 - 9 are zoom in/out and up/down, respectively. The “static” types are omitted in (b) and (c), which stand for fixed distances.
Table 2: Comparison of FMC with other methods. FMC excels in controlling 6D poses of objects and camera with diverse motion patterns.
@@
3DTrajMaster [10] Object Pose Limited Patterns 360°360°360  °-Motion [10]
@@
Figure 5: Motion trajectories. The examples of horizontal and non-horizontal motions indicate the variety and complexity of the trajectories generated by our rule-based algorithm.
$\bullet$ Generation of Multi-Object Scenes. Objects from the same environment with comparable sizes and speeds are selected. The first object’s trajectory is created using the way described earlier, while subsequent trajectories are derived based on the preceding object’s path with a reasonable offset. For camera motion, we randomly track an object in each motion segment, with the trajectory generated using the same method as previously outlined.
$\bullet$ Rendering. This stage creates synthetic videos based on selected assets and motion types of objects and the camera. We divide a complete motion trajectory into many small segments. For each segment, we randomly select animations for objects and generate trajectories according to their annotated motion types. Similarly, random combinations of camera motion types are applied across different segments, allowing the generated videos to encompass diverse motion patterns as in real-world scenarios. These segments are seamlessly combined to form a complex and diverse global trajectory, as shown in Fig. 5. Finally, the selected assets and poses of the camera and objects are imported into Unreal Engine [39] to get 3D animation sequence for rendering. Fig. 1 shows an example trajectory segment.